OpenStudy (anonymous):

Nine times a number minus three is equal to five times that same number added to seven. What is the value of the number?

OpenStudy (uri):

9 times a number ,lets suppose the number is x therefore, 9x-3=5x+7 Now solve for x and get your answer.

OpenStudy (hba):

@pattyadams21 Let me teach you how you can do what @uri has shown on your own :)

OpenStudy (hba):

@pattyadams21 Do you understand what @uri has shown ?

OpenStudy (anonymous):

Yes. I do understand. Surprisingly I wrote the same equation too. Just wanted to make sure it was somewhat right.

OpenStudy (hba):

@pattyadams21 Good job . \[\Huge{\bf{WELCOME \ TO \ OS}}\] Here is your medal for solving it on your own and nailing it right.

OpenStudy (mayankdevnani):

uri's left answer- 9x-3=5x+7 4x=10 x=5/2

OpenStudy (mayankdevnani):

ok @pattyadams21

OpenStudy (hba):

@mayankdevnani I think she can solve that dude.

OpenStudy (mayankdevnani):

may be

OpenStudy (hba):

An interaction is necessary,there is no learning when you just give away the answer.

OpenStudy (anonymous):

It's fine. I just wanted to make sure I wrote the equation right.

OpenStudy (hba):

Yeah you did :D

OpenStudy (anonymous):

Yay me. (:

OpenStudy (hba):

Haha :) , If you are done please close your question :) ,You can also give a best response to the best answer.

OpenStudy (anonymous):

Alrighty then.